# Interactive Inference: A Neuromorphic Theory of Human-Computer Interaction

### v3

@@ 6.1. Study Literature Review
In-depth analysis of crash data shows that nearly one third of all fatal and injury crashes involve driver inattention (Wundersitz, 2019). Lee et al. identified a 30% (310 ms) increase in reaction time when a speech-based email system is used, with implications for safety (Lee et al., 2001). Strayer et al. demonstrated that inattentional blindness can be caused by concurrent cell phone usage while driving, significantly increasing the risk of an accident (Strayer et al., 2011). In a study involving adult drivers, Strayer and Drews observed an 18% slower reaction time while conversing on a cell phone (Strayer and Drews, 2004). Systematic reviews of distracted driving, such as those by Simmons et al. (Simmons et al., 2016, 2017) and Caird et al. (Caird et al., 2014) overwhelmingly demonstrate that tasks that require the drivers’ visual attention, such as dialing, texting, and in-car infotainment touch screen use generate more risk than tasks that do not require visual attention, as these direct the driver’s attentive resources away from the road and thus the primary task (Niu et al., 2019). Researchers have shown particular interest in the car following task due to its prevalence in driving. Wolfe et al’s information acquisition theory modeled the information a driver acquires through peripheral vision and eye movements (Wolfe et al., 2022). Boer pioneered the use of a behavioral entropy measure as an index of workload, and suggested that information entropy (the average amount of information) is the basis for driving behaviour (Boer, 2000). Boer et al. used the sample entropy algorithm to gauge performance of drivers, as demonstrated through a driving simulator study (Boer et al., 2005). Senders did pioneering work on occlusion distance: the distance a driver feels comfortable driving with their eyes shut (Senders et al., 1967). Kujala developed this metric further as a way to measure visual demand while driving (Kujala et al., 2016). Building on Wilde’s risk homeostasis theory (Wilde, 1982), Lu et al. developed a desired safety margin model as a quantified index of risk perception when car following (Lu et al., 2013). That model is based on the leading and following car speeds, and the relative distance between them. Zhu et al. analyzed the driver’s choice of headway (the time interval between two cars at a given speed) in various traffic conditions using a large database (Zhu et al., 2016). Wu et al. proposed a risk repulsion factor F for car-following as determined by the time headway, relative speed, and safe spatial headway (Wu et al., 2021). The driver’s cognitive load has also been assessed and captured as a metric when car-following (Summala, 2000). This is important especially when secondary tasks consume some or all of that load. Subjective questionnaires are often carried out to collect the participants’ assessment of task difficulty, risk, comfort and effort. For example, Lewis-Evans et al. identified a negative relationship between the time headway in a car following task and ratings of task difficulty in a simulator experiment (Lewis-Evans et al., 2010). However, there are many issues with subjective assessments, including but not limited to bias caused by participants’ preferences towards socially desirable responses (Lajunen and Summala, 2003). Others have proposed various physiological measurements as indicators of the subject’s mental effort, such as heart rate variability, electromyogram, and facial temperature (Mulder, 1992; de Waard, 1996; Kajiwara, 2014). Yet such portable physiological sensing solution are unreliable, while trajectory-based risk assessment methods are too complex (Petelczyc et al., 2020).
6.2. KL Divergence of Overlap
@@
Figure 3. Two cars with normal distributions $G(x)$ and $P(x)$ representing uncertainty in their position. The percentage of overlap between distributions equals the probability of a collision (in red). The negative logarithm of this probability gives the number of bits of difference.
[…]
[One car follows the other along a horizontal axis.]One car follows the other along a horizontal axis, labelled as P(x) and Q(x), with a distance of A. The front car’s location has a variation of W along the horizontal axis.
@@
$H$ is the amount of Bayesian surprise, in bits, that the driver would need to process in a car following task. However, recall that the capacity C of the driver’s ability to process this information is only logarithmic. We can establish parameters $a$ and $b$ of this logarithm by measuring the amount of information $H$ using Equation 30 and by performing a regression on the result (we discuss how we operationalized this in Section 7.1):
@@
6.3. Task
We designed a controlled car following task in a driving simulator to collect driving data in target avoidance conditions with varying noise-to-signal ratios. Our goal with this experiment was to model the drivers’ information load using our equations to gauge the amount of information used to maintain a distance $S$ when following a lead car with variation $N$ (see Figure 3). We focused our experimental analysis on deceleration events that may lead to collision.
$S$ and $N$ were varied between conditions. As in real driving, $N$ was not static within a condition: the lead car accelerated and decelerated to different speeds, which added a normally distributed noise $N$ to each distance condition. This allowed us to capture participants’ handling of the car in a range of driving difficulty levels defined by the constituting $N/S$ ratios. We hypothesized that driver’s information capacity during decelerations would be logarithmic with the noise-to-signal ratio, with a mean capacity $b$, in bits of information per bit of difficulty. Because driver information capacity is logarithmic, but the amount of information generated is quadratic, at capacity $b$, the amount of information will outstrip the driver’s ability to process. We hypothesized that this will lead to an error statistic that follows the KL divergence in Equation 30.
6.4. Participants
@@
6.5. Experimental Setup
We presented a simulated driving task with simulated road conditions using BeamNG (BeamNG, 2013), a car simulator video game chosen for its realism, physics engine, low cost, and programmability. Each participant driver was seated in a simulator in front of a curved 49-inch ultrawide monitor with a 32:9 screen ratio. The monitor was positioned 0.6 m from the participant at eye level to provide an immersive simulator view of a car dashboard and windscreen (see Figure 4). A Logitech G29 steering wheel and pedal unit with three pedals were mounted under the simulator rig. Participants were asked to adjust their seat position so that the steering wheel and accelerator/brake pedals were at a comfortable distance. The pedal clutch was not used. Every participant used the same custom car model with the same physics characteristics in BeamNG.
Figure 4. Participant’s experimental setup. The white bar in the center of the screen helped participants perceive the correct distance $S$. They needed to match the size of this bar with the lead car’s bumper while driving. Here, it is wider than the lead car’s bumper indicating that the participant needs to accelerate.
[…]
[Two hands hold a simulator steering wheel in front of a monitor.]Two hands hold a simulator steering wheel in front of a monitor, which displays an in-car field-of-view, a white bar in the middle, a speedometer on the right, and a leading car’s rear bumper. The white bar is between the leading car and the user’s car.
@@ 6.5. Experimental Setup
6.6. Experiment Design
We used a within-subject experiment where the task was to follow the lead car at a specified distance $S$. We then varied the speed of the lead car with variation $N$. Each participant experienced 12 different task conditions with two factors $S$ and $N$.
@@ 6.6. Experiment Design
To constitute each distribution, for each level of $N$, the lead car randomly decelerated 20 times and randomly accelerated 20 times, although always keeping within the level’s range of speed values. Each trial within a level of $N$ comprised a new speed adjustment, which occurred approximately every two seconds. We randomized the presentation order of conditions. We also randomized the presentation order of acceleration or deceleration trials, with accelerations never occurring within a deceleration trial or vice versa.
We measured the instantaneous noise-to-signal ratio (change in distance by lead car $\Delta S$, i.e., some sample of $N$; mean distance to lead car $\overline{S}$, per trial), and pedal activity by the driver. As will be detailed further, our dependent variables were the amount of information processed during a trial (KL divergence), and error (collisions). Together, the 40 trials created a condition simulating a continuous driving task of 2 minutes.
6.7. Procedure
@@ 7. Results
7.1. Measurement and Data Analysis
In our experiment, lead car accelerations never resulted in a collision. Given our interest in the relationship between target avoidance and the risk of collision (error), we constrained our analysis to trials where participants responded to a deceleration of the lead car by braking or by releasing the accelerator pedal. We used 3 different levels of precision for noise-to-signal measurements in our analysis. Each used a different confidence interval for defining $N$.
@@ 7.1. Measurement and Data Analysis
    At the trial level of precision, $N$ was defined by the actual relative backwards distance traveled by the lead car during each trial; and
@@
    Within each trial, $N$ was defined by the standard deviation $(\sigma)$ of a one-second time window of instantaneous distance measurements, as detailed below.
7.1.1. Empirical Factors Noise (N) and Signal (S)
@@
7.1.2. Empirical Noise-to-Signal Ratio
As each trial involved a braking event that resulted in processing of information given an individually administered $\Delta S$, many similar $N/S$ ratios were created across conditions. We therefore used these more detailed empirical observations of noise-to-signal ratios for our log-linear regression analysis, pooled across all trials. Here, our definition of noise-to-signal ratio was $\Delta S$ per trial divided by the measured mean $S$ ($\overline{S}$) per trial, yielding the metric $\Delta S/\overline{S}$. We also used this definition for our regression analysis of error rates, as this yielded directly comparable results of error rates vs. capacity, in bits. This definition was also used for our Index of Difficulty (see Equation 34). Note that the only effect of the use of sampled $\Delta S$ values instead of some confidence interval $N$ is a scaling effect on $N/S$ ratios.
7.1.3. Measuring Information Processed
Our first dependent variable was the amount of information processed by the driver during decelerations. To measure the instantaneous amount of information processed within a trial, we used a more detailed 1-second rolling time window to sample the distribution of distances $S$ to the lead car at 450 Hz. Our definition of noise $N$ here was the standard deviation ($\sigma$) of the distribution of $S$ obtained in this time window, while our definition of signal $S$ was the mean $S$ ($\overline{S}$) of this time window. First, the noise-to-signal ratio was sampled using this type of one second window at the beginning of each trial, ending at the moment the driver responded to the stimulus. We then calculated the amount of information in the stimulus by entering this first noise-to-signal ratio into Equation 30. A one-second window again sampled the noise-to-signal ratio at the end of each trial, prior to (i.e., ending at) a new deceleration or acceleration event. We calculated the remaining amount of information by entering this second noise-to-signal ratio into Equation 30.
Subtracting the remaining amount of information from the amount of information in the stimulus gave us the amount of information processed by the driver during each deceleration trial. We subsequently aggregated measurements across all conditions, sorting results using the $\Delta S/\overline{S}$ from Section 7.1.2, and binning measurements that were within 0.05 units of $\Delta S/\overline{S}$. While this means our definitions for signal-to-noise were different between y and x axes, we note that our instantaneous KL measurements of information (y-axis) are orthogonal to and independent from the definition of empirical signal-to-noise ratio per trial used for the index of difficulty.
7.1.4. Measuring Error
@@
7.2. Effects of Factors N and S
Table 1 shows the mean KL divergence values for consumed information against levels of factors $N$ and $S$. Since the residuals in our ANOVA model of the dependent variable differed significantly from a normal distribution (KS test, p¿0.05), we used a non-parametric Friedman test. The Friedman test showed highly significant effects of both factors $N$ ($\chi^{2}(2)=17.73,p<0.001$) and $S$ ($\chi^{2}(3)=40.84,p<0.001$) on the number of bits consumed by the driver (note that a two-way ANOVA produced the same results).
7.3. Car Following Model
Table 1. Mean KL (s.e.) for consumed information (in bits) for levels of factors $N$ and $S$.
[…]
[A table of consumed information in bits of 4 levels S against 3 levels N.]A table of consumed information in bits of 4 levels S per column (2.84/4.84/6.84/8.84) against 3 levels N per row (1.93/4.19/7.20). Consumed bits range from 0.01, of N 8.84 m x S 1.93 m, to 0.17, of W 2.84 m x A 7.20 m. S N 2.84 m 4.84 m 6.84 m 8.84 m 1.93 m 0.11 (0.07) 0.06 (0.04) 0.03 (0.02) 0.01 (0.02) 4.19 m 0.16 (0.09) 0.09 (0.04) 0.05 (0.03) 0.04 (0.04) 7.20 m 0.17 (0.09) 0.10 (0.03) 0.08 (0.04) 0.05 (0.02)
Figure 5. Driver information capacity plotted against noise-to-signal ratio ($\Delta S/\overline{S}$, index of difficulty (ID) on a logarithmic scale below x-axis). Capacity followed a logarithmic function (blue solid), with a slope of 0.25 bits (green dash). Black solid line indicates remaining capacity.
[…]
[A curve plot of three lines, and a series of data dots along the rising logarithmic line.]A curve plot of three lines: 1) Line 1 shows driver information capacity following a logarithmic function, 2) Line 2 shows a slope of 0.25 bits which is parallel to x axis, and 3) Line 3 shows a dropping line indicating remaining capacity. Line 1 and Line 2 intersects around 1.2 bit, and Line 1 and Line 3 intersects around 0.5 bit. The x-axis is noise-to-signal ratio on a logarithmic scale. The data dots spread along Line 1, the driver information capacity.
@@ 7.3. Car Following Model
The mean capacity (bandwidth $b$) of our car following task was measured at $b=0.25$ bits/bit, while the Index of Performance (throughput $1/b$) was 3.96 bits/bit. The y axis indicates bits of consumed information, while the x axis indicates noise-to-signal ratio $\Delta$S/$\overline{S}$, or bits of $ID$ if the bottom log scale is used. Note that the only difference in plotting Figure 5 on a log-linear scale against $ID$, as is customary when plotting with Fitts’ law, is that lines would be straight, with the bottom scale as the primary x-axis. Mean capacity is expressed as bits of information consumed per bit of difficulty, while the Index of Performance (1/b) is expressed as the bits of difficulty processed per bit of consumed information.
7.4. Error (Collision) Model
Figure 6. Error (non-surprise, in bits) plotted against noise-to-signal ratio ($\Delta S/\overline{S}$, index of difficulty (ID) on a logarithmic scale below x-axis). Error followed a KL divergence (red dots). Blue solid lines shows driver capacity and black solid line remaining capacity.
@@ 7.4. Error (Collision) Model
[A curve plot of three lines, and a series of data dots along the rising power line.]A curve plot of three lines: 1) Line 1 shows an error curve following a KL divergence with data dots along it, 2) Line 2 shows a driver capacity line following a logarithmic curve, and 3) Line 3 shows a dropping remaining capacity curve. Line 1 and Line 2 intersects around 1.6 bit, and Line 1 and Line 3 intersects around 0.7 bit. The x-axis is noise-to-signal ratio on a logarithmic scale.
@@
Figure 6 shows the relationship between error (red solid) and aggregated signal-to-noise measurements on a $\Delta S/\overline{S}$ scale across conditions, as well as on a logarithmic scale ($ID$, bottom), compared with driver capacity (blue solid) and remaining capacity (black solid). The total number of collisions across all trials was 90. Error rates below or equal to a noise-to-signal ratio or $ID$ of 1 trended, on average, slightly below the expected 4% (at 0.05 bits or 3.4%, between 2 and 4 collisions across trials), while those above 1 increased quadratically, averaging 0.38 bits (23%), with a maximum of 0.82 bits (43% or 39 collisions across trials) at $ID=1.85$.
7.5. Qualitative Results
Figure 7. Perceived task difficulty for each condition: participants’ responses to the question “This driving task was difficult” on a scale of 1 (Strongly Disagree) to 7 (Strongly Agree).
[…]
[A bar chart of all conditions’ subjective ratings aligned vertically by the middle-point choice of “Neither Agree nor Disagree”.]A bar chart of all conditions’ subjective ratings (7-point scale) aligned vertically by the middle-point (level-4) choice of “Neither Agree nor Disagree”. In each row, the choices are sorted by the rating from 1 (“Strongly Disagree”) to 7 (“Strongly Agree”), left to right. For the easiest condition (row 1), the range of the user’s rating is -0.80–0.13. For the hardest condition (row 12), the range of the user’s rating is 0.00–1.00.
Figure 7 shows participants’ questionnaire responses ranking difficulty per conditional $N/S$, sorted by $ID$. These represent mean responses to the question “This driving task was difficult”, on a scale from 1 (Strongly Disagree) to 7 (Strongly Agree) (Lewis-Evans et al., 2010). A Friedman non-parametric analysis of variance showed participants rated the car following task as significantly more difficult with increasing $ID$ ($\chi^{2}(11)=116.4,p<0.0001$). Note that all participants found the task to be easy to interpret and execute, while none reported distraction by the floating bar graphic during the task.
8. Discussion

### v4

@@ Introduction
In-depth analysis of crash data shows that nearly one third of all fatal and injury crashes involve driver inattention (Wundersitz, 2019). Lee et al. identified a 30% (310 ms) increase in reaction time when a speech-based email system is used, with implications for safety (Lee et al., 2001). Strayer et al. demonstrated that inattentional blindness can be caused by concurrent cell phone usage while driving, significantly increasing the risk of an accident (Strayer et al., 2011). In a study involving adult drivers, Strayer and Drews observed an 18% slower reaction time while conversing on a cell phone (Strayer and Drews, 2004). Systematic reviews of distracted driving, such as those by Simmons et al. (Simmons et al., 2016, 2017) and Caird et al. (Caird et al., 2014) overwhelmingly demonstrate that tasks that require the drivers’ visual attention, such as dialing, texting, and in-car infotainment touch screen use generate more risk than tasks that do not require visual attention, as these direct the driver’s attentive resources away from the road and thus the primary task (Niu et al., 2019). Researchers have shown particular interest in the car following task due to its prevalence in driving. Wolfe et al.’s information acquisition theory modeled the information a driver acquires through peripheral vision and eye movements (Wolfe et al., 2022). Boer pioneered the use of a behavioral entropy measure as an index of workload, and suggested that information entropy (the average amount of information) is the basis for driving behaviour (Boer, 2000). Boer et al. used the sample entropy algorithm to gauge performance of drivers, as demonstrated through a driving simulator study (Boer et al., 2005). Senders did pioneering work on occlusion distance: the distance a driver feels comfortable driving with their eyes shut (Senders et al., 1967). Kujala developed this metric further as a way to measure visual demand while driving (Kujala et al., 2016). Building on Wilde’s risk homeostasis theory (Wilde, 1982), Lu et al. developed a desired safety margin model as a quantified index of risk perception when car following (Lu et al., 2013). That model is based on the leading and following car speeds, and the relative distance between them. Zhu et al. analyzed the driver’s choice of headway (the time interval between two cars at a given speed) in various traffic conditions using a large database (Zhu et al., 2016). Wu et al. proposed a risk repulsion factor F for car-following as determined by the time headway, relative speed, and safe spatial headway (Wu et al., 2021). The driver’s cognitive load has also been assessed and captured as a metric when car-following (Summala, 2000). This is important especially when secondary tasks consume some or all of that load. Subjective questionnaires are often carried out to collect the participants’ assessment of task difficulty, risk, comfort and effort. For example, Lewis-Evans et al. identified a negative relationship between the time headway in a car following task and ratings of task difficulty in a simulator experiment (Lewis-Evans et al., 2010). However, there are many issues with subjective assessments, including but not limited to bias caused by participants’ preferences towards socially desirable responses (Lajunen and Summala, 2003). Others have proposed various physiological measurements as indicators of the subject’s mental effort, such as heart rate variability, electromyogram, and facial temperature (Mulder, 1992; de Waard, 1996; Kajiwara, 2014). Yet such portable physiological sensing solution are unreliable, while trajectory-based risk assessment methods are too complex (Petelczyc et al., 2020).
KL Divergence of Overlap
@@
Figure 3. Two cars with normal distributions $G(x)$ and $P(x)$ representing uncertainty in their position. The percentage of overlap between distributions equals the probability of a collision. The negative logarithm of this probability gives the number of bits of difference.: One car follows the other along a horizontal axis.
@@
$H$ is the amount of Bayesian surprise, in bits, that the driver would need to process in a car following task. However, recall that the capacity C of the driver’s ability to process this information is only logarithmic. We can establish parameters $a$ and $b$ of this logarithm by measuring the amount of information $H$ using Equation 32 and by performing a regression on the result (we discuss how we operationalized this in Section Measurement and Data Analysis):
@@
Task
We designed a controlled car following task in a driving simulator to collect driving data in target avoidance conditions with varying noise-to-signal ratios. Our goal with this experiment was to model the drivers’ information load using our equations to gauge the amount of information used to maintain a distance $S$ when following a lead car with variation NN (see Figure 3). We focused our experimental analysis on deceleration events that may lead to collision.
$S$ and NN were varied between conditions. As in real driving, NN was not static within a condition: the lead car accelerated and decelerated to different speeds, which added a normally distributed noise NN to each distance condition. This allowed us to capture participants’ handling of the car in a range of driving difficulty levels defined by the constituting $N/S$ ratios. We hypothesized that driver’s information capacity during decelerations would be logarithmic with the noise-to-signal ratio, with a mean capacity $b$, in bits of information per bit of difficulty. Because driver information capacity is logarithmic, but the amount of information generated is quadratic, at capacity $b$, the amount of information will outstrip the driver’s ability to process. We hypothesized that this will lead to an error statistic that follows the KL divergence in Equation 32.
Participants
@@
Experimental Setup
We presented a simulated driving task with simulated road conditions using BeamNG (BeamNG, 2013), a car simulator video game chosen for its realism, physics engine, low cost, and programmability. Each participant driver was seated in a simulator in front of a curved 49-inch monitor with a 32:9 screen ratio. The monitor was positioned 0.6 m from the participant at eye level to provide an immersive simulator view of a car dashboard and windscreen (see Figure 4). A Logitech G29 steering wheel and pedal unit with three pedals were mounted under the simulator rig. Participants were asked to adjust their seat position so that the steering wheel and accelerator/brake pedals were at a comfortable distance. The pedal clutch was not used. Every participant used the same custom car model with the same physics characteristics in BeamNG.
Figure 4. Participant’s experimental setup. The white bar in the center of the screen helped participants perceive the correct distance $S$. They needed to match the size of this bar with the lead car’s bumper while driving. Here, it is wider than the lead car’s bumper indicating that the participant needs to accelerate.: Two hands hold a simulator steering wheel in front of a monitor.
@@
Experiment Design
We used a within-subject experiment where the task was to follow the lead car at a specified distance $S$. We then varied the speed of the lead car with variation NN. Each participant experienced 12 different task conditions with two factors $S$ and NN.
@@
To constitute each distribution, for each level of NN, the lead car randomly decelerated 20 times and randomly accelerated 20 times, although always keeping within the level’s range of speed values. Each trial within a level of NN comprised a new speed adjustment, which occurred approximately every two seconds. We randomized the presentation order of conditions. We also randomized the presentation order of acceleration or deceleration trials, with accelerations never occurring within a deceleration trial or vice versa.
We measured the instantaneous noise-to-signal ratio (change in distance by lead car $\Delta S$, i.e., some sample of NN; mean distance to lead car $\overline{S}$, per trial), and pedal activity by the driver. As will be detailed further, our dependent variables were the amount of information processed during a trial (KL divergence), and error (collisions). Together, the 40 trials created a condition simulating a continuous driving task of 2 minutes.
Procedure
@@ Results
Measurement and Data Analysis
In our experiment, lead car accelerations never resulted in a collision. Given our interest in the relationship between target avoidance and the risk of collision (error), we constrained our analysis to trials where participants responded to a deceleration of the lead car by braking or by releasing the accelerator pedal. We used 3 different levels of precision for noise-to-signal measurements in our analysis. Each used a different confidence interval for defining NN.
@@
    At the trial level of precision, NN was defined by the actual relative backwards distance traveled by the lead car during each trial; and
@@
    Within each trial, NN was defined by the standard deviation $(\sigma)$ of a one-second time window of instantaneous distance measurements, as detailed below.
Empirical Factors Noise (N) and Signal (S)
@@
Empirical Noise-to-Signal Ratio
As each trial involved a braking event that resulted in processing of information given an individually administered $\Delta S$, many similar $N/S$ ratios were created across conditions. We therefore used these more detailed empirical observations of noise-to-signal ratios for our log-linear regression analysis, pooled across all trials. Here, our definition of noise-to-signal ratio was $\Delta S$ per trial divided by the measured mean $S$ ($\overline{S}$) per trial, yielding the metric $\Delta S/\overline{S}$. We also used this definition for our regression analysis of error rates, as this yielded directly comparable results of error rates vs. capacity, in bits. This definition was also used for our Index of Difficulty (see Equation 36). Note that the only effect of the use of sampled $\Delta S$ values instead of some confidence interval NN is a scaling effect on $N/S$ ratios.
Measuring Information Processed
Our first dependent variable was the amount of information processed by the driver during decelerations. To measure the instantaneous amount of information processed within a trial, we used a more detailed 1-second rolling time window to sample the distribution of distances $S$ to the lead car at 450 Hz. Our definition of noise NN here was the standard deviation ($\sigma$) of the distribution of $S$ obtained in this time window, while our definition of signal $S$ was the mean $S$ ($\overline{S}$) of this time window. First, the noise-to-signal ratio was sampled using this type of one second window at the beginning of each trial, ending at the moment the driver responded to the stimulus. We then calculated the amount of information in the stimulus by entering this first noise-to-signal ratio into Equation 32. A one-second window again sampled the noise-to-signal ratio at the end of each trial, prior to (i.e., ending at) a new deceleration or acceleration event. We calculated the remaining amount of information by entering this second noise-to-signal ratio into Equation 32.
Subtracting the remaining amount of information from the amount of information in the stimulus gave us the amount of information processed by the driver during each deceleration trial. We subsequently aggregated measurements across all conditions, sorting results using the $\Delta S/\overline{S}$ from Section Empirical Noise-to-Signal Ratio, and binning measurements that were within 0.05 units of $\Delta S/\overline{S}$. While this means our definitions for signal-to-noise were different between y and x axes, we note that our instantaneous KL measurements of information (y-axis) are orthogonal to and independent from the definition of empirical signal-to-noise ratio per trial used for the index of difficulty.
Measuring Error
@@
Effects of Factors N and S
Table 1 shows the mean KL divergence values for consumed information against levels of factors NN and $S$. Since the residuals in our ANOVA model of the dependent variable differed significantly from a normal distribution (KS test, p¿0.05), we used a non-parametric Friedman test. The Friedman test showed highly significant effects of both factors NN ($\chi^{2}(2)=17.73,p<0.001$) and $S$ ($\chi^{2}(3)=40.84,p<0.001$) on the number of bits consumed by the driver (note that a two-way ANOVA produced the same results).
Car Following Model
Table 1. Mean KL (s.e.) for consumed information (in bits) for levels of factors NN and $S$.: A table of consumed information in bits of 4 levels S against 3 levels N.
S
N 2.84 m 4.84 m 6.84 m 8.84 m
1.93 m 0.11 (0.07) 0.06 (0.04) 0.03 (0.02) 0.01 (0.02)
4.19 m 0.16 (0.09) 0.09 (0.04) 0.05 (0.03) 0.04 (0.04)
7.20 m 0.17 (0.09) 0.10 (0.03) 0.08 (0.04) 0.05 (0.02)
Figure 5. Driver information capacity plotted against noise-to-signal ratio ($\Delta S/\overline{S}$, index of difficulty (ID) on a logarithmic scale below x-axis). Capacity followed a logarithmic function (blue solid), with a slope of 0.25 bits (green dash). Black solid line indicates remaining capacity.: A curve plot of three lines, and a series of data dots along the rising logarithmic line.
@@
Error (Collision) Model
Figure 6. Error (non-surprise, in bits) plotted against noise-to-signal ratio ($\Delta S/\overline{S}$, index of difficulty (ID) on a logarithmic scale below x-axis). Error followed a KL divergence (red dots). Blue solid lines shows driver capacity and black solid line remaining capacity.: A curve plot of three lines, and a series of data dots along the rising power line.
@@
Figure 6 shows the relationship between error (red solid) and aggregated signal-to-noise measurements on a $\Delta S/\overline{S}$ scale across conditions, as well as on a logarithmic scale (IDID, bottom), compared with driver capacity (blue solid) and remaining capacity (black solid). The total number of collisions across all trials was 90. Error rates below or equal to a noise-to-signal ratio or IDID of 1 trended, on average, slightly below the expected 4% (at 0.05 bits or 3.4%, between 2 and 4 collisions across trials), while those above 1 increased quadratically, averaging 0.38 bits (23%), with a maximum of 0.82 bits (43% or 39 collisions across trials) at $ID=1.85$.
Qualitative Results
Figure 7. Perceived task difficulty for each condition: participants’ responses to the question “This driving task was difficult” on a scale of 1 (Strongly Disagree) to 7 (Strongly Agree).: A bar chart of all conditions’ subjective ratings aligned vertically by the middle-point choice of “Neither Agree nor Disagree”.
Figure 7 shows participants’ questionnaire responses ranking difficulty per conditional $N/S$, sorted by IDID. These represent mean responses to the question “This driving task was difficult”, on a scale from 1 (Strongly Disagree) to 7 (Strongly Agree) (Lewis-Evans et al., 2010). A Friedman non-parametric analysis of variance showed participants rated the car following task as significantly more difficult with increasing IDID ($\chi^{2}(11)=116.4,p<0.0001$). Note that all participants found the task to be easy to interpret and execute, while none reported distraction by the floating bar graphic during the task.
Discussion